OpenStudy (anonymous):

A spherical fish bowl is half-filled with water. The center of the bowl is C, and the length of segment AB is 20 inches, as shown below. Use Twenty two over seven for pi. A sphere with diameter 20 inches is drawn. Which of the following can be used to calculate the volume of water inside the fish bowl?

OpenStudy (anonymous):

1 over 24 over 322 over 7(203) 1 over 24 over 322 over 7(202)(10) 1 over 24 over 322 over 7(103) 1 over 24 over 322 over 7(102)(20) @calculusxy

OpenStudy (calculusxy):

\[V = \frac{ 4 }{ 3 }\pi r^2\]

OpenStudy (calculusxy):

Our diameter is 20 inches, and radius is always half of diameter. Thus, the radius is 20/2 which equals .... ?

OpenStudy (anonymous):

r=10

OpenStudy (calculusxy):

yes now plug those info into the formula given

OpenStudy (anonymous):

so V=4 over 3 pi10^2??

OpenStudy (calculusxy):

yes \[\frac{ 4 }{ 3 }\pi (10)^2\] \[\frac{ 4 }{ 3 }(\frac{ 22 }{ 7 })10^2\]

OpenStudy (calculusxy):

Remember that pi in here means 22/7

OpenStudy (anonymous):

disregard that comment

OpenStudy (anonymous):

But that isn't any of the options given

OpenStudy (calculusxy):

This looks like a complex question especially with the multiple choices.

OpenStudy (calculusxy):

@Nnesha

OpenStudy (calculusxy):

Hopefully Nnesha can help you... I am sorry

OpenStudy (anonymous):

@calculusxy lol it's fine at least you tried